# From Peptides to Nanostructures: A Euclidean Transformer for Fast and Stable Machine Learned Force Fields

### v1

@@ II.3 Data Efficiency, Stability and Extrapolation
Figure 6: (a) Dissociation of the AT-AT complex over time, due to the breaking of $\pi$-$\pi$ interactions. (b) Velocity auto-correlation function for the buckyball catcher (upper) and the double-walled nanotube (lower). For the nanotube, the structure is shown from the side (i) and from the front (ii). (c) Temperature dependency of the velocity auto-correlation function, investigated along the DHA molecule for three different temperatures. All auto-correlation functions have been obtained from MD simulations over 1 ns.
@@
A subtle case is highlighted by the adenine-thymine complex (AT-AT). The MD simulations show one instability (in a total of six runs) for the equivariant model at 500 K, which illustrates that the stability improvement of an equivariant model should be considered as a reduction of the chance of failure rather than a guarantee for stability. We remark that unexpected behaviors can not be ruled out for any empirical model. We further observed dissociation of substructures (either A, T or AT) from the AT-AT complex during MD simulations (Fig. 6 (a.ii)). Such a behavior corresponds to the breaking of hydrogen bonds or $\pi$-$\pi$-interactions, which highlights weak interactions as a challenge for MLFFs. Interestingly, for other supra-molecular structures the non-covalent interactions are described correctly (section II.4 and Fig. 1 (b)). The training data for AT-AT has been sampled from a 20 ps long ab-initio MD trajectory which only covers a small subset of all possible conformations and makes it likely to leave the data manifold. As a consequence, we observe an increase in the rate of dissociation when increasing the simulation temperature, since it effectively extends the space of accessible conformations per unit simulation time.
II.4 From Peptides to Nanostructures
Velocity auto-correlation functions are an important tool to relate MD simulations to real world experimental data. Here, we calculate velocity auto-correlation functions for systems ranging from small peptides up to host-guest systems and nanostructures. To achieve a correct description for such systems, the model must describe non-covalent bonding correctly and be stable for nanoseconds of simulation time. For the largest structure with 370 atoms, 5M MD steps with SO3krates takes 20h simulation time ($\sim 15$ ms per step).
We train an individual model for each structure in the MD22 data set and compare it to the sGDML model (Tab. 3). To that end, we decided to train the model on two different sets of training data sizes: (A) On structure depended sizes (600 to 8k) as reported in [55], and (B) on structure independent sizes of 1k training points per structure. Since some settings might require accurate predictions when trained on a smaller number of training data points, we chose to include setting (B) into our analysis. The approximation accuracies achievable with SO3krates compare favourably to the ones that have been observed with the sGDML model [32, 55] (Tab. 3). Even for setting (B) the force errors on the test set are below 1 kcal molÅ${}^{-1}\,$\mathrm{\SIUnitSymbolAngstrom}$^{-1}$. We use the SO3krates FFs to run 1 ns long MD simulations, which enables the calculation of converged velocity auto-correlation functions and a comparison to experimental data from IR spectroscopy. We start by analysing two supra-molecular structures in form of a host-guest system and a small nanomaterial. The former play an important role for a wide range of systems in chemistry and biology [56, 25], whereas the latter offer promises for the design of materials with so far unprecedented properties [57]. Here, we investigate the applicability of the SO3krates FF to such structures on the example of the buckyball catcher and the double walled nanotube (Fig. 6 (b)).
@@ II.4 From Peptides to Nanostructures
For DHA, we further analyze the evolution of the velocity auto-correlation function with temperature and find non-trivial shifts in the spectrum hinting towards the capability of the model to learn non-harmonic contributions of the PES. As pointed out in [59], FFs that only rely on (learn) harmonic bond and angle approximations fail to predict changing population or temperature shifts in the middle to high frequency regime. Similar results are obtained for Ac-Ala3-NHMe (SI Fig. 11).
II.5 Potential Energy Surface Topology
@@ III Discussion
Long time-scale MD simulations are essential to reveal converged dynamic and thermodynamic observables of molecular systems [62, 63, 64, 65]. Despite achieving low test errors, many state-of-the-art MLFFs exhibit unpredictable behavior caused by the accumulation of unphysical contributions to the output, making it extremely difficult or even impossible to reach extended timescales [30]. This prevents the extraction of physically faithful observables at scale. Ongoing research aims at improving stability by incorporating physically meaningful inductive biases via various kinds of symmetry constraints [8, 11, 17, 66, 33, 34, 67], but the large computational cost of current solutions mitigates many practical advantages.
@@
Applying the SO3krates architecture to different structures from the MD22 benchmark, including peptides (Ac-Ala3-NHMe, DHA) and supra-molecular structures (AT-AT, buckyball catcher, double walled nanotube), yields stable molecular dynamics (MD) simulations with impressive time scales of tens of nanoseconds per day. This enables the computation of converged velocity auto-correlation functions, allowing comparison to experimental measurements. We have also shown, that SO3krates reliably reveals conformational changes in small bio-molecules on the example of DHA and Ac-Ala3-NHMe. To that end, SO3krates is able to predict physically valid minima conformations which have not been part of the training data. The representative nature of Ac-Ala3-NHMe holds the potential that a similar behavior can be obtained for much larger peptides and proteins. The limited availability of ab-initio data for structures at this scale, makes extrapolation to unknown parts of the PES a crucial ingredient on the way to large scale bio-molecular modeling.
While our development makes stable extended simulation timescales accessible using modern MLFF modeling paradigms in an unprecedented manner, future work remains to be done in order to bring the applicability of MLFFs even closer to that of conventional classical FFs. Various encouraging avenues in that direction are currently emerging: In the current design, the EV are only defined in terms of two-body interactions. Recent results suggest that accuracy can be further improved by incorporating atomic cluster expansions into the MP step [68, 69, 47]. At the same time, this may help reducing the number of MP steps which in turn decreases the computational complexity of the model. Another, yet open discussion is the appropriate treatment of global effects. Promising steps have been taken by using low-rank approximations [24, 70], trainable Ewald summation [71] or by adding long-range corrections from continuum solvent theory [72]. Further, a recent work showed that adding long-range interactions can improve the accuracy on the MD22 benchmark [73].
Future work will therefore focus on the seamless incorporation of many-body expansions, global effects, and long-range interactions into the EV formalism and aim to further increase computational efficiency to ultimately bridge MD time-scales at high accuracy.
@@ IV.5 MD Stability
We define an MD to be stable when (A) there is an uncontrolled dissociation of the system, and (B) each bond length follows a reasonable distribution over time. We refer to failure mode (A) as an explosion of the MD simulation when (at least one of) the force predictions of the MLFF diverges during the MD simulation (Fig. 8 (a). A decomposition of (parts of) the molecule can be detected by a strong peak in MD temperature, which is usually a few orders of magnitude larger than the target temperature. We assume a bond length to be distributed reasonably, when it does not differ by more than 50 % from the equilibrium bond distance at any point of the simulation. Criteria (A) has e.g. been used in [30] to determine the MD stability of different MLFFs. However, in certain cases analysing MD temperature can be an insufficient condition to detect unstable behavior, e.g. when single bonds dissociate slowly over time or take on non-physical values over a temporarily limited interval (Fig. 8 (b)). Such a behavior, however, is easily identified using criteria (B).
@@ IV.8 Network and Training
All other models use degrees $l=\{1,2,3\}$ in the EV, $T=3$ and initialize the EV according to Eq. (11). For the MD17 stability experiments, 10,000 conformations are randomly selected of which 9,500 are used for training and 500 for validation. For the MD22 benchmark a varying number of training samples plus 500 validation samples or 1000 training samples plus 500 validation samples are drawn randomly. The models trained on 1000 samples are used for the calculation of the velocity auto-correlation functions and for the minima hopping experiments.
@@ A.1 MD Simulations
In Fig. 9 and Fig. 10 we plot the distribution of total energy values, energies as a function of time and the temperature as a function of time for the MD simulations which have been used to calculate the velocity auto-correlation functions (main body Fig. 6). As it can be readily verified, the performed simulations are energy conserving where the total energies follow a Gaussian distribution with a small, but non-zero variance that is a consequence of the finite step size in the numerical integration performed in the Velocity-Verlet update.
Figure 11 (b) shows the velocity auto-correlation function for Ac-Ala3-NHMe at three different temperatures of 100 K, 300 K and 500 K. As reported for DHA in the main body of the text, we find non-trivial shifts in frequency and population across different temperatures. On the right hand side, we show the radial and angular distribution functions for both Ac-Ala3-NHMe and DHA at a temperature of 500 K.
@@ A.3 DHA Data Efficiency
Figure 9: Total energy distribution, total energy over time (inset) and the temperature over time as observed in the MD simulations for DHA with target temperatures (a) 100 K, (b) 300 K and (c) 500 K using the Velocity-Verlet algorithm. From the resulting trajectories, the velocity auto-correlation function reported in the main body of the text have been calculated.
@@
Figure 11: Velocity auto-correlation function for Ac-Ala3-NHMe at different temperatures obtained with the SO3krates-FF.

### v2

@@ II.3 Data Efficiency, Stability and Extrapolation
Figure 6: Distribution of the radius of gyration (ROG) for Ac-Ala3-NHMe as a function of MD simulation time in 20 ns steps. The distribution converges after 60-80 ns simulation time, underlining the importance of stable, but at the same time computationally efficient, simulations. The small inset at the top shows the ROG in Angstrom as a function of simulation time.
A subtle case is highlighted by the adenine-thymine complex (AT-AT). The MD simulations show one instability (in a total of six runs) for the equivariant model at 500 K, which illustrates that the stability improvement of an equivariant model should be considered as a reduction of the chance of failure rather than a guarantee for stability. We remark that unexpected behaviors can not be ruled out for any empirical model. We further observed dissociation of substructures (either A, T or AT) from the AT-AT complex during MD simulations (Fig. 8 (a.ii)). Such a behavior corresponds to the breaking of hydrogen bonds or $\pi$-$\pi$-interactions, which highlights weak interactions as a challenge for MLFFs. Interestingly, for other supra-molecular structures the non-covalent interactions are described correctly (section II.6 and Fig. 1 (b)). The training data for AT-AT has been sampled from a 20 ps long ab-initio MD trajectory which only covers a small subset of all possible conformations and makes it likely to leave the data manifold. As a consequence, we observe an increase in the rate of dissociation when increasing the simulation temperature, since it effectively extends the space of accessible conformations per unit simulation time.
II.4 Radius of Gyration
The radius of gyration (ROG) is an important observable for determining the structural and dynamical behavior of polymers as it allows to gain an estimate for structural compactness of proteins [62] and is experimentally accessible [63]. The time scales of structural changes are often between tenths or even hundreds of nanoseconds [64] which requires simulation durations at the same order of magnitude to observe a converged distribution of the ROG. Thus, the MLFF must be robust enough to yield stable dynamics for hundreds of nanoseconds while being computationally efficient at the same time.
Figure 7: Dynamics of Ala15 obtained from a SO3krates model, trained on only 1k data points of a much smaller peptide (Ac-Ala3-NHMe). Analysis of the end-to-end difference shows rapid folding into helical structure illustrating the generalization capabilities of the learned local representations towards conformational changes on length-scales greatly exceeding the training data (dashed grey line).
Here we showcase the potential of the proposed SO3krates model for such applications by using it to calculate a converged distribution of the ROG for Ac-Ala3-NHMe from 100 ns long MD simulations at 300 K (Fig. 6). With a time step of $0.5$ fs such a simulation requires 200M force evaluations. The SO3krates model enables us to perform such simulations within 5 days on a single A100 GPU. By analysing the ROG distribution as a function of simulation time (different colors in Fig. 6) we find such time scales to be necessary for convergence. We find characteristic peaks in the distribution, which correspond to the folded and unfolded conformation, respectively. Details on the MD simulation can be found in the methods section.
II.5 Generalization to Larger Peptides
Generalization to larger structures and unknown conformations is an inevitable requirement for scaling MLFFs to realistic simulations in biochemistry. Here, we showcase that by only using 1k training points of a small peptide (42 atoms), SO3krates can generalize to much larger peptides (151 atoms) without the need of any additional training data. Despite the locality of the model, we observe folding into a helical structure illustrating the extrapolation capabilities of the learned representations to larger structures.
We use the same model as for the ROG experiment from section II.4. As already illustrated, the obtained model is able to perform long and stable dynamics for the structure it has been trained on. To further increase the complexity of the task, we use the model without any modification to investigate the dynamics of Ala15, starting from the extended structure (most left structure in Fig. 7). Our analysis of the end-to-end distance between the carbonyl carbon of the first residue and the last residue (green spheres in Fig. 7) reveals that the peptide rapidly folds into the secondary, helical structure (most right structure in Fig. 7). A comparison to the end-to-end distance in Ac-Ala3-NHMe (dashed horizontal line) reveals the generalization capabilities of SO3krates towards conformational changes on length-scales that go beyond the ones present in the training data (grey dashed line in Fig. 7).
II.6 Power Spectra
Power spectra of the atom velocities are an important tool to relate MD simulations to real world experimental data. They are calculated as the Fourier transform of the velocity auto-correlation function for systems ranging from small peptides up to host-guest systems and nanostructures. To achieve a correct description for such systems, the model must describe both covalent and non-covalent bonding correctly. For the largest structure with 370 atoms, 5M MD steps with SO3krates takes 20h simulation time ($\sim 15$ ms per step).
Figure 8: (a) Dissociation of the AT-AT complex over time, due to the breaking of $\pi$-$\pi$ interactions. (b) Power spectra for the buckyball catcher (upper panel) and the double-walled nanotube (lower panel) computed as the Fourier transform of the velocity auto-correlation function. For the nanotube, the structure is shown from the side (i) and from the front (ii). (c) Power spectrum for DHA from an NVE at the zero point energy (light blue) and 500 K (dark blue), as well as the frequencies from harmonic approximation (red dashed lines).
@@ II.6 Power Spectra
We train an individual model for each structure in the MD22 data set and compare it to the sGDML model (Tab. 3). A comparison of training time to other neural network architectures is given in the methods section IV.9. To that end, we decided to train the model on two different sets of training data sizes: (A) On structure depended sizes (600 to 8k) as reported in [65], and (B) on structure independent sizes of 1k training points per structure. Since some settings might require accurate predictions when trained on a smaller number of training data points, we chose to include setting (B) into our analysis. The approximation accuracies achievable with SO3krates compare favourably to the ones that have been observed with the sGDML model [35, 65] (Tab. 3). Even for setting (B) the force errors on the test set are below 1 kcal molÅ${}^{-1}\,$\mathrm{\SIUnitSymbolAngstrom}$^{-1}$. We use the SO3krates FFs to run 1 ns long MD simulations, from which we calculate the power spectra , enabling a comparison to experimental data from IR spectroscopy. Although the frequencies in these systems do not require such simulation lengths, we chose them to illustrate computational feasibility as well as simulation stability. We start by analysing two supra-molecular structures in form of a host-guest system and a small nanomaterial. The former play an important role for a wide range of systems in chemistry and biology [66, 26], whereas the latter offer promises for the design of materials with so far unprecedented properties [67]. Here, we investigate the applicability of the SO3krates FF to such structures on the example of the buckyball catcher and the double walled nanotube (Fig. 8 (b)).
@@
For DHA, we further analyse the evolution of the velocity auto-correlation function with temperature and find non-trivial shifts in the spectrum hinting towards the capability of the model to learn non-harmonic contributions of the PES. As pointed out in [69], FFs that only rely on (learn) harmonic bond and angle approximations fail to predict changing population or temperature shifts in the middle to high frequency regime. To further showcase the anharmonicity of the spectra obtained with SO3krates, we first identify the global minimum of DHA (using the minima hopping results from section II.7). For the found conformation, we calculate harmonic frequencies as well as the zero point energy (ZPE) from harmonic approximation. The ZPE is 12.979 eV, which corresponds to a temperature of roughly 930 K. We find that our method yields stable dynamics, even though the simulation temperature is almost twice as high as the one used for generating the training data (500 K). The emergence of non-trivial shifts between the spectra from the NVE (blue curve) with the harmonic frequencies (dashed red lines) illustrates the non-trivial anharmonicity that our proposed method is able to model. Similar results are obtained for Ac-Ala3-NHMe (SI Fig. 13).
II.7 Potential Energy Surface Topology
@@ III Discussion
Long time-scale MD simulations are essential to reveal converged dynamic and thermodynamic observables of molecular systems [72, 73, 64, 74, 75]. Despite achieving low test errors, many state-of-the-art MLFFs exhibit unpredictable behavior caused by the accumulation of unphysical contributions to the output, making it extremely difficult or even impossible to reach extended timescales [32]. This prevents the extraction of physically faithful observables at scale. Ongoing research aims at improving stability by incorporating physically meaningful inductive biases via various kinds of symmetry constraints [8, 11, 17, 76, 36, 37, 77], but the large computational cost of current solutions mitigates many practical advantages.
@@
Applying the SO3krates architecture to different structures from the MD22 benchmark, including peptides (Ac-Ala3-NHMe, DHA) and supra-molecular structures (AT-AT, buckyball catcher, double walled nanotube), yields stable molecular dynamics (MD) simulations with remarkable time scales on the order of tens of nanoseconds per day. This enables the computation of experimentally relevant observables, including power spectra and converged distributions for the radius of gyration for small peptides. We have also shown, that SO3krates reliably reveals conformational changes in small bio-molecules on the example of DHA and Ac-Ala3-NHMe. To that end, SO3krates is able to predict physically valid minima conformations which have not been part of the training data. The representative nature of Ac-Ala3-NHMe holds the potential that a similar behavior can be obtained for much larger peptides and proteins. The limited availability of ab-initio data for structures at this scale, makes extrapolation to unknown parts of the PES a crucial ingredient on the way to large scale bio-molecular modeling.
In a recent line of work, methods have been proposed that aim to reduce the computational complexity of SO(3) convolutions [78, 79]. They serve as a drop-in replacement for full SO(3) convolutions whereas the method presented here allows to fully avoid expensive SO(3) convolutions within the message passing paradigm. This result as well as ours demonstrate that the optimization of equivariant interaction is an active research field that has not yet fully matured, potentially offering further paths for improvement.
While our development makes stable extended simulation timescales accessible using modern MLFF modeling paradigms in an unprecedented manner, future work remains to be done in order to bring the applicability of MLFFs even closer to that of conventional classical FFs. Various encouraging avenues in that direction are currently emerging: In the current design, the EV are only defined in terms of two-body interactions. Recent results suggest that accuracy can be further improved by incorporating atomic cluster expansions into the MP step [80, 81, 52, 82]. At the same time, this may help reducing the number of MP steps which in turn decreases the computational complexity of the model.
Another, yet open discussion is the appropriate treatment of global effects. Promising steps have been taken by using low-rank approximations [24, 83], trainable Ewald summation [84] or by learning long-range corrections in a physically inspired way [85, 86, 87, 88, 89, 17, 24, 90, 91]. Approaches of the latter type are of particular importance when extrapolation to larger systems is required. Although equivariant models can improve the extrapolation capabilities for local interactions, this does not hold for interactions that go beyond the length scales present in the training data or beyond the effective cutoff of the model. Since the aforementioned methods rely on local properties such as partial charges [17, 24, 87, 88], electronegativities [89] or Hirshfeld volumes [85, 86] they can be seamlessly integrated into our method by learning a corresponding local descriptor in the invariant feature branch of the SO3krates architecture.
Future work will therefore focus on the incorporation of many-body expansions, global effects, and long-range interactions into the EV formalism and aim to further increase computational efficiency to ultimately bridge MD time-scales at high accuracy.
@@ IV.5 MD Stability
We define an MD simulation to be stable when (A) there is no non-physical dissociation of the system, and (B) each bond length follows a reasonable distribution over time. We refer to failure mode (A) when there is a divergence in the predicted velocities resulting in instantaneous dissociation of all or some atoms in the molecule which corresponds to non-physical behavior (Fig. 10 (a)). In contrast, a physically meaningful dissociation happens on reasonable time and kinetic energy scales. A decomposition of (parts of) the molecule can be detected by a strong peak in MD temperature, which is usually a few orders of magnitude larger than the target temperature. We assume a bond length to be distributed reasonably, when it does not differ by more than 50 % from the equilibrium bond distance at any point of the simulation. Criteria (A) has e.g. been used in [32] to determine the MD stability of different MLFFs. However, in certain cases analysing MD temperature can be an insufficient condition to detect unstable behavior, e.g. when single bonds dissociate slowly over time or take on non-physical values over a temporarily limited interval (Fig. 10 (b)). Such a behavior, however, is easily identified using criteria (B).
@@ IV.8 Network and Training
All other models use degrees $l=\{1,2,3\}$ in the EV, $T=3$ and initialize the EV according to Eq. (11). For the MD17 stability experiments, 10,000 conformations are randomly selected of which 9,500 are used for training and 500 for validation. For the MD22 benchmark a varying number of training samples plus 500 validation samples or 1000 training samples plus 500 validation samples are drawn randomly. The models trained on 1000 samples are used for the calculation of the spectra , for the minima hopping experiments and for the radius of gyration analysis.
@@
IV.9 Training Step Time
The total training time can strongly vary depending on selection of hyperparameters such as batch size or the number of epochs until convergence of the loss, which itself depends on the number of data points, the learning rate and the optimizer that is used. To make a comparison as fair as possible it thus makes sense to report the time per gradient update when comparing MLFF models.
In Ref. [95], the training time consumption has been compared for different state-of-the-art models, where e.g. PaiNN [37] takes $\sim 10$ ms, VISNet [95] takes $\sim 50$ ms and NequIP[36] or Allegro [53] take $\sim 200$ ms for Chignolin (166 atoms). When training on Chignolin we achieve a training time consumption of $\sim 8$ ms measured on an A100GPU. The training times reported in  [95] are measured on a V100 GPU, which is a factor 1.6 to 3.4 slower than an A100, depending on the data modality and architecture. Thus, our runtime comparison has an uncertainty of the same factor, but can illustrate an order of magnitude faster training step speed compared to models such as NequIP or Allegro.
The gradient step time depends on the inference speed of the model as well as on the number of parameters in the model. As already shown SO3krates outperforms other equivariant models that are competitive in stability and accuracy in the terms of inference speed. Further, SO3krates has only 311k parameters which is lightweight compared to other equivariant models as ALLEGRO or NequIP. Thus, our measured runtimes for training are also in line with our expectation from theoretical considerations.
@@ A.1 MD Simulations
In Fig. 11 and Fig. 12 we plot the distribution of total energy values, energies as a function of time and the temperature as a function of time for the MD simulations which have been used to calculate the power spectra (main body Fig. 8). As it can be readily verified, the performed simulations are energy conserving where the total energies follow a Gaussian distribution with a small, but non-zero variance that is a consequence of the finite step size in the numerical integration performed in the Velocity-Verlet update.
Figure 13 (b) shows the power spectra for Ac-Ala3-NHMe at three different temperatures of 100 K, 300 K and 500 K. As reported for DHA in the main body of the text, we find non-trivial shifts in frequency and population across different temperatures. On the right hand side, we show the radial and angular distribution functions for both Ac-Ala3-NHMe and DHA at a temperature of 500 K.
@@ A.3 DHA Data Efficiency
Figure 11: Total energy distribution, total energy over time (inset) and the temperature over time as observed in the MD simulations for DHA with target temperatures (a) 100 K, (b) 300 K and (c) 500 K using the Velocity-Verlet algorithm. From the resulting trajectories, the power spectra reported in the main body of the text have been calculated.
@@
Figure 13: Power spectra for Ac-Ala3-NHMe at different temperatures obtained with the SO3krates-FF.